# Dipole response in deformed halo nuclei $^{42}\mathrm{Mg}$ and $^{44}\mathrm{Mg}$

### v1

@@ I Introduction
During the past decades, the relativistic density functional theory (DFT) has been proven to be powerful in successfully describing various nuclear properties and predicting new phenomena Meng (2016). In order to describe the exotic nuclei, by taking into account the pairing correlations and continuum effects in a microscopic and self-consistent way, the relativistic continuum Hartree-Bogoliubov (RCHB) theory was developed Meng and Ring (1996); Meng (1998). The RCHB theory successfully describes the ground states of the spherical nuclei both near and far from the $\beta$-stability line Meng et al. (2006), interprets the halo in $^{11}\mathrm{Li}$ Meng and Ring (1996), and predicts the giant halo in Zr Meng and Ring (1998) and Ca Meng et al. (2002) isotopes. To describe the deformed exotic nuclei, the deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) was developed by solving the RHB equations in a Dirac Woods-Saxon basis Zhou et al. (2010); Li et al. (2012). Inheriting the advantages of the RCHB theory and including the deformation degrees of freedom, the DRHBc theory has been applied to study deformed exotic nuclei, including $^{17,19}\mathrm{B}$ Yang et al. (2021); Sun (2021a), $^{22}\mathrm{C}$ Sun et al. (2018), $^{31}\mathrm{Ne}$ Zhong et al. (2022); Pan et al. (2024), $^{37}\mathrm{Mg}$ Zhang et al. (2023a), and to predict the halos in $^{39}\mathrm{Na}$ Zhang et al. (2023b) and $^{42,44}\mathrm{Mg}$ Zhou et al. (2010); Li et al. (2012); Zhang et al. (2019); Sun and Zhou (2021).
@@ II.1 Deformed relativistic Hartree-Bogoliubov theory in continuum
The detailed formalism of the DRHBc theory can be seen in Refs. Li et al. (2012); Zhang et al. (2020). In the DRHBc theory, the mean field and pairing correlations are treated self-consistently by the relativistic Hartree-Bogoliubov (RHB) equation,
@@ IV Results and discussion
Fig. 3 shows the single-neutron levels around the Fermi surface in the canonical basis for $^{40,42,44}\mathrm{Mg}$. Following Refs. Zhou et al. (2010); Li et al. (2012), the single-neutron levels in $^{42}\mathrm{Mg}$ and $^{44}\mathrm{Mg}$ can be naturally divided into two parts: the well-bound levels with $\varepsilon<-2\ \mathrm{MeV}$ corresponding to the “core” part, and the remaining levels with $\varepsilon>-1\ \mathrm{MeV}$ corresponding to the “halo” part. To understand the microscopic structures of the $K^{\pi}=1^{-}$ states below 3 MeV in $^{42}\mathrm{Mg}$ and $^{44}\mathrm{Mg}$, the QRPA amplitudes in the quasiparticle basis $X_{\mu\nu}^{i}$ and $Y_{\mu\nu}^{i}$ of these states are extracted from the QFAM amplitudes $X_{\mu\nu}(\omega)$ and $Y_{\mu\nu}(\omega)$ using the method from Ref. Hinohara et al. (2013). These amplitudes are then transformed into the canonical basis to get the amplitudes $X_{ab}^{i}$ and $Y_{ab}^{i}$ with a,ba,\,b denoting the canonical states. After the summation of the indices $b$, the quantity $\sum_{b}(|X_{ab}^{i}|^{2}-|Y_{ab}^{i}|^{2})$ measures the contribution of transitions from the single-particle level labeled by $a$. Within this framework, the dominant transition for each state can be determined, and is schematically labeled in Fig. 3. It is found that the state at 2.50 MeV in $^{42}\mathrm{Mg}$ is predominantly (59.7%) generated by such canonical states $ab$: the index $a$ corresponds to the $1/2^{-}$ single-neutron level at 0.44-0.44 MeV, and $b$ corresponds to various single-neutron levels with $\varepsilon>0\ \mathrm{MeV}$. In other words, this specific state is dominated by transitions from the $1/2^{-}$ single-neutron level. Similarly, the state at 2.56 MeV in $^{42}\mathrm{Mg}$ is dominated by transitions from the $3/2^{-}$ single-neutron level at 0.22-0.22 MeV (65.2%); the state at 2.30 MeV in $^{44}\mathrm{Mg}$ is dominated by transitions from the $1/2^{-}$ single-neutron level at 0.91-0.91 MeV (59.5%); the state at 2.39 MeV in $^{44}\mathrm{Mg}$ is dominated by transitions from the $3/2^{-}$ single-neutron level at 0.86-0.86 MeV (68.5%). In short, the $K^{\pi}=1^{-}$ states below 3 MeV in $^{42}\mathrm{Mg}$ and $^{44}\mathrm{Mg}$ are dominated by transitions from the “halo” part of the single-neutron levels.

### v2

@@ I Introduction
During the past decades, the relativistic density functional theory (DFT) has been proven to be powerful in successfully describing various nuclear properties and predicting new phenomena Meng (2016). In order to describe the exotic nuclei, by taking into account the pairing correlations and continuum effects in a microscopic and self-consistent way, the relativistic continuum Hartree-Bogoliubov (RCHB) theory was developed Meng and Ring (1996); Meng (1998). The RCHB theory successfully describes the ground states of the spherical nuclei both near and far from the $\beta$-stability line Meng et al. (2006), interprets the halo in $^{11}\mathrm{Li}$ Meng and Ring (1996), and predicts the giant halo in Zr Meng and Ring (1998) and Ca Meng et al. (2002) isotopes. To describe the deformed exotic nuclei, the deformed relativistic Hartree-Bogoliubov theory in continuum (DRHBc) was developed by solving the relativistic Hartree-Bogoliubov (RHB) equations in a Dirac Woods-Saxon basis Zhou et al. (2010); Li et al. (2012). Inheriting the advantages of the RCHB theory and including the deformation degrees of freedom, the DRHBc theory has been applied to study deformed exotic nuclei, including $^{17,19}\mathrm{B}$ Yang et al. (2021); Sun (2021a), $^{22}\mathrm{C}$ Sun et al. (2018), $^{31}\mathrm{Ne}$ Zhong et al. (2022); Pan et al. (2024), $^{37}\mathrm{Mg}$ Zhang et al. (2023a), and to predict the halos in $^{39}\mathrm{Na}$ Zhang et al. (2023b) and $^{42,44}\mathrm{Mg}$ Zhou et al. (2010); Li et al. (2012); Zhang et al. (2019); Sun and Zhou (2021).
@@ II.1 Deformed relativistic Hartree-Bogoliubov theory in continuum
The detailed formalism of the DRHBc theory can be seen in Refs. Li et al. (2012); Zhang et al. (2020). In the DRHBc theory, the mean field and pairing correlations are treated self-consistently by the RHB equation,
@@ IV Results and discussion
Fig. 3 shows the single-neutron levels around the Fermi surface in the canonical basis for $^{40,42,44}\mathrm{Mg}$. In $^{42}\mathrm{Mg}$, the gap between the $5/2^{-}$ level at 2.45-2.45 MeV and the $1/2^{-}$ level at 0.44-0.44 MeV is $2.01$ MeV. In $^{44}\mathrm{Mg}$, the gap between the $5/2^{-}$ level at 2.85-2.85 MeV and the $1/2^{-}$ level at 0.91-0.91 MeV is $1.94$ MeV. Following Refs. Zhou et al. (2010); Li et al. (2012), the single-neutron levels in $^{42}\mathrm{Mg}$ and $^{44}\mathrm{Mg}$ can be naturally divided into two parts: the well-bound $5/2^{-}$ level and the below ones corresponding to the “core” part, and the remaining levels corresponding to the “halo” part. To understand the microscopic structures of the $K^{\pi}=1^{-}$ states below 3 MeV in $^{42}\mathrm{Mg}$ and $^{44}\mathrm{Mg}$, the QRPA amplitudes in the quasiparticle basis $X_{\mu\nu}^{i}$ and $Y_{\mu\nu}^{i}$ of these states are extracted from the QFAM amplitudes $X_{\mu\nu}(\omega)$ and $Y_{\mu\nu}(\omega)$ using the method from Ref. Hinohara et al. (2013). These amplitudes are then transformed into the canonical basis to get the amplitudes $X_{ab}^{i}$ and $Y_{ab}^{i}$ with a,ba,\,b denoting the canonical states. After the summation of the indices $b$, the quantity $\sum_{b}(|X_{ab}^{i}|^{2}-|Y_{ab}^{i}|^{2})$ measures the contribution of transitions from the single-particle level labeled by $a$. Within this framework, the dominant transition for each state can be determined, and is schematically labeled in Fig. 3. It is found that the state at 2.50 MeV in $^{42}\mathrm{Mg}$ is predominantly (59.7%) generated by such canonical states $ab$: the index $a$ corresponds to the $1/2^{-}$ single-neutron level at 0.44-0.44 MeV, and $b$ corresponds to various single-neutron levels with $\varepsilon>0\ \mathrm{MeV}$. In other words, this specific state is dominated by transitions from the $1/2^{-}$ single-neutron level. Similarly, the state at 2.56 MeV in $^{42}\mathrm{Mg}$ is dominated by transitions from the $3/2^{-}$ single-neutron level at 0.22-0.22 MeV (65.2%); the state at 2.30 MeV in $^{44}\mathrm{Mg}$ is dominated by transitions from the $1/2^{-}$ single-neutron level at 0.91-0.91 MeV (59.5%); the state at 2.39 MeV in $^{44}\mathrm{Mg}$ is dominated by transitions from the $3/2^{-}$ single-neutron level at 0.86-0.86 MeV (68.5%). In short, the $K^{\pi}=1^{-}$ states below 3 MeV in $^{42}\mathrm{Mg}$ and $^{44}\mathrm{Mg}$ are dominated by transitions from the “halo” part of the single-neutron levels.